OpenStudy (anonymous):

when asked to factor the trinomial x^2-14x+49, a student gives the answer (x-7) (x+7). Which of the following statements is true? A. The answer is correct B. The answer is incorrect; the 7s should be 14s C. The answer is incorrect; the minus sign should be a plus sign D. The answer is incorrect; the plus sign should be a minus sign.

jhonyy9 (jhonyy9):

so you know that x^2 -14x +49 you can rewriting it like x*x - 7x -7x +(-7)(-7) yes ? so from first two elements you can factorizing out x and will get x(x-7) and from 3rd and 4thd element you can factorizing out 7 and will get -7(x-7) so tha you rewrite this equation like x^2 -14x +49=x(x-7)-7(x-7) but here you can factorizing out (x-7) and will get x^2 -14x +49 = (x-7)(x-7) so x^2 -14x +49 = (x-7)^2 so from these result that D. is right sure hope so much that these is understandably sure right easy

OpenStudy (denisaboichuk):

d